For the year ended December 31, 2018, Carla Co. reported pretax financial income of $228,176. Its current tax payable was $59,445. Carla reported a difference between pretax financial statement income and taxable income. This difference is due to accelerated depreciation for income tax purposes. Carla’s income tax rate is 29% and Carla made no estimated tax payments during 2018.

What amount of accelerated depreciation did take in 2018?
Answer #1

Taxable income = 59445/29% = 204,983 (rounded off)

Accelerated depreciation = 228,176 - 204,983 = 23,193

amount of accelerated depreciation in 2018 = 23,193

Any doubt please comment, If it helped you, please give a thumbs up.